A car of mass 2500 kg travels at a velocity of 75 m/s due west. What is the magnitude and direction of the momentum of the car


1
Expert's answer
2021-06-21T16:02:43-0400
"p=mv\\p=(2500)(75)=187500\frac{kg\cdot m}{s}"

 Direction: due west.